OpenStudy (anonymous):

What is the slope–intercept form of the equation 4x − 2y = 9? A. y =2x +9/2 B.y =9x +2/9 C. y =9x −2/9 D. y =2x −9/2

OpenStudy (sidsiddhartha):

slope intercept form is y=mx+b so 4x-2y=9 2y=4x-9 y=2x-9/2

OpenStudy (sidsiddhartha):

do u get it?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

oh and mr 2 jews in the ove how about naw